OpenStudy (anonymous):

Solve the following system. y = x + 4 5x + y = 16 Answer (2, -6) (2, 6) (6, 2) (-2, 6)

OpenStudy (anonymous):

y = x + 4 so 5x + x + 4 = 16 6x + 4 = 16 x = 12/6 x = 2 so, y = 2 + 4 y = 6 therefore, (2,6)

OpenStudy (anonymous):

as (x,y)